?help differential eq by syste elimnation?
Solve the given system of differential equations by systematic elimination. dx = -x + 2 dt dy = y + z dt dz = -x + y dt I
Answer #1

We solve for y take y= emt - y=memt y = me m² emt gam3cmt Hence in (4) we have (m²+2 m² + m). emt = m +2m² +mao ( ent & o) 7+ Cl₂ + (3) e-tt Hence, the general solution to the system of equations, y (t) = Citlert & Catert z (t) = CitCzert where C.,